# OpenUAS: Embeddings of Cities in Japan with Anchor Data for Cross-city Analysis of Area Usage Patterns

### v2

@@ Abstract
We publicly release OpenUAS, a dataset of area embeddings based on urban usage patterns, including embeddings for over 1.3 million 50-meter square meshes covering a total area of 3,300 square kilometers. This dataset is valuable for analyzing area functions in fields such as market analysis, urban planning, transportation infrastructure, and infection prediction. It captures the characteristics of each area in the city, such as office districts and residential areas, by employing an area embedding technique that utilizes location information typically obtained by GPS. Numerous area embedding techniques have been proposed, and while the public release of such embedding datasets is technically feasible, it has not been realized. One of the obstacles has been the integration of data from different cities and periods into a unified space without sharing raw location data. We address this issue by developing an anchoring method that establishes anchors within a shared embedding space. We publicly release this anchor dataset along with area embedding datasets from several periods in eight major Japanese cities. This dataset allows users to analyze urban usage patterns in Japanese cities and embed their urban dataset into the same embedding space using the anchoring method. Our key contributions include the development of the anchoring method, releasing area embedding datasets for Japanese cities, and providing tools for effective data utilization.
@@
On the other hand, area embedding techniques have been proposed, aggregating features derived from location information by area and mapping them onto a vector space[29, 30, 31, 32, 33]. Such embeddings allow for advanced area analysis by aggregating information for each area, which eliminates personal data and enables searching and clustering of areas with similar functions and usage patterns. In this paper, we apply the embedding method called Area2Vec proposed by Shoji et al[34]. This method embeds areas based on stay information extracted from GPS data, resulting in embeddings called UAS (Usage of Area with Stay information) that reflect how people use the areas. The UAS can capture not only the static functions of urban areas but also the dynamic changes in area functions by tracking shifts in stay patterns due to environmental changes(urban development, seasonal changes, pandemics). By adopting UAS as the public area embedding dataset, it becomes possible to analyze changes in area functions over multiple periods, in addition to comparing area functions between cities.
Releasing area embedding datasets provides valuable information for analyzing area functions to users who cannot directly access raw GPS datasets. Despite the demand for the public release of area embedding datasets, existing methods have not released area embedding datasets. One reason is that area embeddings, learned independently in each dataset, are unique representations. They embed the relative similarities between areas within a dataset into an embedding space. This uniqueness poses a challenge: areas from different datasets, such as cities or periods, cannot be directly compared. For example, embeddings learned independently for areas in New York and Tokyo cannot be compared due to the lack of a common embedding space. Similarly, embeddings of the same city from different periods, such as pre-COVID and post-COVID periods, are incomparable. Therefore, we cannot compare them with existing publicly available embeddings, even if new embeddings are generated for a different city (or a different period) using their datasets. This issue can be addressed by combining datasets of New York and Tokyo or by training new translation models[35, 36]. However, if the location data for each city is owned by different organizations, sharing the actual location data becomes difficult, making these methods impractical.
@@
  • The public release of area embedding data for eight major cities in Japan with anchor data based on extensive location information.
  • Development and release of tools and codes for effective use of area embedding data.
First, the next section will detail the area embedding and anchoring methods. In Data Records, we will explain the specifics of the area embedding data, and in Validation, we will assess the effectiveness and anonymity of the embedding representations and the anchoring method. Finally, in Usage Notes, we will discuss effective ways to utilize area embedding data, including the anchoring method, and provide relevant code.
@@ Methods
First, we will explain the GPS dataset we use for area embeddings and the dataset released in this paper. We release area embeddings using the Area2Vec method and the anchoring method described later. Area2Vec embeds areas based on stay information in each area. We first obtained GPS data, which includes anonymized user IDs, latitude, longitude, and timestamps, from Blogwatcher Inc. This data is obtained with the use of multiple applications on mobile devices with user consent. To extract stay information from the GPS data, we used the stay estimation method by Iwata et al[37]. The resulting stay dataset includes columns for latitude, longitude, day of the week, time, and stay duration for each stay, where the day of the week and time indicate the start of the stay.
@@
The architecture of Area2Vec is shown in Figure 1[34]. The input is a one-hot vector corresponding to each area ID in the dataset, and the output is a one-hot vector corresponding to each discretized stay feature. As shown in Table 3, the discretization of stay features involves assigning the stay day of the week, arrival time(start time of each stay), and stay duration to discrete classes. This discretization results in the stay characteristics being classified into $2*12*7=168$ categories, and therefore, in this paper, $N^{\prime}=168$. The weight of the input layer W corresponds to the Area Embedding, and by learning the relationships between areas and discretized stay features, areas with similar stay features are embedded close to each other in the vector space, while dissimilar areas are embedded further apart. In more detail, when a one-hot vector of each area ID is inputted, the model is trained to approximate a 168-dimensional statistical vector, which represents the aggregated discretized stay characteristics of that area. Consequently, areas with similar statistical vectors will have alike vectors, resulting in the 168 dimensions being compressed into a $D_{area}$ dimensional embedding. Therefore, clustering of area embeddings yields results as shown in Figure 4. This figure will be explained in detail in the Validation section.
Figure 1: Area2Vec architecture (Adapted from Shoji et al., 2024[34]).
Day of week Weekday or Holiday
@@
We develop an anchoring method to fix the embedding space across datasets, allowing the embedding of areas within each dataset into a common space. This enables the comparison of area embeddings across different datasets. First, we sample area data (anchor data) that includes typical areas such as residential and office districts. Then, during the embedding process, this anchor data is mixed with the dataset for training, fixing the embedding values $W_{A}$ corresponding to the Anchor. This approach fixes the latent meanings of each dimension in the embedding space across datasets. The following sections will sequentially explain the details of the anchoring method.
Figure 2: Anchor data is sampled from location data. Data from multiple cities and periods are embedded and clustered. Samples from each cluster are used as anchor data, and reference values are obtained as embeddings for each anchor.
First, we sample a group of anchor data $D_{a}=\{d_{a_{n}}\}$, including typical area data such as residential and office districts. The overview of the sampling of anchor data is shown in Figure 2. Each anchor data $d_{a_{n}}$ corresponds to the $n$th anchor $a_{n}$. Each anchor $a_{n}$ corresponds to a landmark vector in the latent space that needs to be evenly distributed. In other words, each anchor data $d_{a_{n}}$ should ideally represent a different area pattern, and $D_{a}$ as a whole should encompass a diverse range of area tendencies. To sample such $D_{a}$, we embedded a large location dataset $D^{*}$ into a single latent space, followed by clustering, and used data sampled from each cluster as the respective anchor data $d_{a_{n}}$. $D^{*}$ includes areas from eight cities in Japan, with data from Nagoya covering the years 2019-2023 (treating the same area as different for each year). The resulting $D_{a}$ includes unique area tendencies for each of the eight cities, as well as a variety of area tendencies such as pre-COVID, during the pandemic, and post-pandemic. For the public release data, we set the number of anchors $n=512$ and the amount of data for each anchor data at $20,000$. Optimization of the number of anchors and the amount of data is described in the appendix. In the $D_{a}$, timestamp and stay duration are recorded in 15-minute increments to ensure that the stays of specific individuals cannot be extracted from $D_{a}$.
@@
The OpenUAS can be accessed at zenodo(https://zenodo.org/records/13141800). Embedding data for each dataset is aggregated every 50m or 250m mesh and stored as vector information. The total number of meshes is 1,418,827. The data columns include geocode, Latitude, Longitude, Geometry, vector, cluster5, cluster10, and cluster20. The geocode is an identifier assigned to each standard mesh, with 250m meshes having 10-digit and 50m meshes having 12-digit integer values. Latitude and Longitude indicate the center coordinates of each mesh. Geometry and Vector are lists of floating-point numbers, with the former representing the mesh polygon and the latter the vector representation of the area embedding. cluster5, cluster10, and cluster20 are the numbers assigned when the area embeddings are clustered into those respective numbers. Graphs showing the stay patterns of each cluster in these clusterings are also included.
Additionally, we released the anchor embeddings and anchor dataset to enable users to embed their own data into the same latent space using the anchoring method. The Anchor Embeddings are defined as 512x8 dimensional tensors in PyTorch and are stored as .pt files. The anchor data is a 512x20,000 record .csv file containing anchor_id, arrival_time, and stay_time. All codes, including those for embedding and area analysis, will be released at GitHub repository(https://github.com/UCLabNU/OpenUAS/). Detailed explanations on how to use them are provided in the Usage Note and README.
Figure 4: Stay patterns of each 5 cluster.
@@
In this section, we demonstrate the utility of our data by practicing a part of cross-city and cross-period area analysis using clustering of area embeddings, thereby showing its effectiveness. Here, we embedded all meshes from all cities and periods into the same latent space and used k-means++ for clustering. The number of clusters was set to five, but this number is arbitrary and can be increased for more detailed analysis. Graphs for when the number of clusters was set to 10 and 20 are included in the appendix section. For different periods of data in Nagoya, we treated each period as a distinct area. For example, the meshes of Nagoya in January 2021 and February 2021 are considered as different meshes. This allows us to analyze how area clusters change over time. The resulting stacked graph is shown in Figure 4. The horizontal axis represents time, and the vertical axis represents the number of visits every 30 minutes, with the color indicating the length of stay. On the right side of the graph, we display labeled interpretations(manually predicted) based on the tendencies of each cluster, along with the count of meshes belonging to each cluster. For instance, the first and second clusters, with many short stays on weekends compared to weekdays, can be inferred as areas used like entertainment districts. The third cluster, with many stays over 360 minutes during weekdays, appears to be an office district cluster, while the fourth cluster, with frequent stays under 30 minutes during weekday mornings and evenings, seems to be areas with stations or streets. The fifth cluster, with many stays over 720 minutes, mainly during weekend nights, is a residential area cluster. Thus, areas can be clustered based on their functional differences, and each interpretation of the cluster can be labeled based on the inferred stay tendencies from the graphs.
Next, we visualize the geographical map of these area clusters. We begin by showing the map of area clusters for each city in April 2021 in Figure 5. From the figure, we can observe the distribution of various areas, such as central train stations and street areas, industrial and office regions in downtown and bay areas, and suburban residential districts. An important point to note here is that despite each area being independently learned, the anchoring method ensures that each color representing an area function has the same meaning across different cities. This allows for comparative analysis, for instance, between the industrial area distributions in Tokyo and Osaka, assessing similarities and differences in industrial zones across cities. Simultaneously, other researchers can embed and analyze areas in other cities using their own data. Thus, our data and method facilitate cross-city analysis of area functions.
@@
Next, we evaluate cross-period area analysis. We present the distribution of area clusters in the central part of Nagoya for April from 2019 to 2023 in Figure 6. Area (a) around Nagoya Station is a region where commercial facilities, restaurants, and offices are located. In 2019, many of these areas belonged to the entertainment and restaurant cluster, but in 2020, many shifted to the workplace cluster. This change reflects the impact of the COVID-19 pandemic in 2020, indicating a decrease in visitors to entertainment and restaurant areas due to travel restrictions. Area (b) is Sakae, a major entertainment area in Nagoya. Here, too, the function of entertainment areas decreased in 2020, but there is a visible recovery of this function in 2021. Area (c) is around Nagoya Castle, a tourist spot, with government buildings such as the city hall to its south. In 2020, this area transitioned to a station and street area, likely due to decreased visitors for entertainment purposes, while visits for commuting to the workplace area south of Nagoya Castle remained. Annual data, as illustrated here, enable analysis of area changes before and after the COVID-19 pandemic.
@@
Finally, we visualize the monthly changes of areas within the same region. We have chosen and visualized the three months from September to November 2021 from the entire year, as shown in Figure7. Area (d) includes a park area with an art museum and a science museum, where a festival event was held from September 26 to October 31, 2021. The impact of this event, coupled with the lifting of the state of emergency from August 27 to September 30, 2021, can be seen in the changing functions of the area. In area (e), a large shopping mall opened on October 27, 2021. This led to the emergence of an entertainment area in a region that previously only had workplaces and residential areas. Thus, using annual and monthly data, it is possible to analyze changes in the functions of the same area over time.
@@
Effectiveness of anchoring method
Next, we validate the effectiveness of the anchoring method. The key aspect of the anchoring method is the ability to embed different datasets into the common embedding space independently. In this method, independent datasets are embedded into a common embedding space $E^{+}$ through the anchor data $D_{a}$. To evaluate whether they are appropriately embedded in the common space, we independently embed $D^{*}$ again. We then define Misalignment, as shown in the following equation, as the discrepancy between the obtained embedding $E$ and $E^{+}$
$Misalignment(E,E^{+})=\frac{1}{|M|}\sum\limits_{m\in M}Dist(e_{m},e^{+}_{m})$ (3)
$M$ represents the set of all meshes in the dataset $D^{*}$, with each mesh denoted as $m$ and the total number of meshes in $M$ as $|M|$. $e_{m}$ represents the retrained embedding corresponding to $m$, whereas $e_{m}^{+}$ represents the embedding for $m$ in $E^{+}$. A smaller $Misalignment(E,E^{+})$ indicates a greater success in embedding each mesh into the same latent space as $E^{+}$ in the retraining. $Dist$ is a distance function for which both Euclidean distance and Cosine distance are used. Cosine distance is calculated as 1 minus the cosine similarity. Here, we compared the vanilla Area2Vec without anchoring and Area2Vec with anchoring. Results are shown in Figure 8. From these results, it is evident that the vanilla method yields different embeddings for independent learning, indicating that it does not embed in a common latent space. In contrast, our proposed anchoring method resulted in a smaller variance in embeddings for the same data, demonstrating the ability to embed areas into the same representation space.
@@
Simultaneously, it is necessary to evaluate whether the performance of area embeddings has been compromised by anchoring. This concern arises because the anchoring method may distort the update of weights during learning via anchors, potentially preventing the acquisition of optimal area embeddings for the training data. In particular, area data with stay patterns not included in anchor data may not be representable in the embedding space of OpenUAS. To validate this, we conducted the following experiments for each dataset comprising $D^{*}$.
@@
If the performance using anchoring significantly deteriorates compared to when learning individually, it may indicate that the anchor data does not fully capture the latent patterns of areas within the target data, failing to anchor them appropriately in the ideal space.
Although area embedding is an unsupervised learning process, poses challenges for direct performance validation. To address this, we introduce the metric $ApproximationLoss$. Area embedding serves as a method of dimension reduction. During the training process, the product of the embedding $e_{m}$ for area m and the output matrix $W^{\prime}$ is optimized to better approximate the frequency $f_{m}$ of the discretized stay feature for each mesh. Here, $f_{m}$ is a 168-dimensional vector representing the frequency of the one-hot for discretized stay features. Essentially, the performance of area embedding improves as the product of $e_{m}$ and $W^{\prime}$ more closely approximates $f_{m}$. This discrepancy is quantified by $ApproximationLoss$. The corresponding formula is presented below.
@@
Here, $f_{m}$ represents the true frequency of discretized stay features in mesh $m$, and $\hat{f_{m}}$ represents the approximate frequency calculated as the product of $e_{m}$ and $W^{\prime}$. We conducted the previously described experiments on $D_{tokyo}$ to $D_{2023}$, which make up $D^{*}$, calculating and comparing the $ApproximationLoss$ for cases learned with and without anchoring. The results are shown in Figure 9. The results reveal that there is no significant difference in the $ApproximationLoss$ values before and after the application of the anchoring method. However, it is also noted that the $ApproximationLoss$ values differ by a few points regardless of the application of the anchoring method, with variations being particularly notable across different years. This variation is believed to be caused by the differences in latent trends of areas across the datasets. For instance, the year 2019 shows worse metrics compared to other datasets, which could be attributed to the greater diversity in area characteristics, making it more challenging to represent them with fewer embedding dimensions. The variations in embedding accuracy due to differences in cities and years suggest that further evaluation is needed using data from cities outside Japan. Nonetheless, the minor deterioration in embedding accuracy before and after applying anchoring indicates that the method is effective.
@@
In the second step, these statistical vectors are embedded into an h-dimensional latent space, further abstracting the information. As shown in the validation section, the degree of match between the approximated and actual statistical vectors for each mesh is calculated by $ApproximationLoss$. Therefore, data users can only know the approximate values of the statistics for each mesh. Moreover, $ApproximationLoss$ tends to be higher for meshes with more user counts and lower for those with fewer. Figure 10 illustrating the relationship between user count per mesh and $ApproximationLoss$ is shown. From Figure 10, it is apparent that in meshes with 10-20 user counts, the values are higher and decrease as the user count increases. This result suggests it is particularly difficult to infer related statistical information in meshes with fewer users. Therefore, our data does not contain personal data and only provides approximate values for any statistical information that could indirectly estimate personal data. From these aspects, it can be concluded that our data is sufficiently anonymized.
@@
This involves clustering a group of target area embeddings into any number of clusters, forming clusters of areas with similar functions. This allows for the classification and analysis of urban sections based on their functions and usage. Our code includes the clustering of area embeddings into any number of clusters and implements the code for geographical visualization.
@@
As explained in the Validation, area trends can be predicted through the Approximation of discretized stay statistics. While not perfectly matching actual statistics, this method can reproduce stay trend graphs as shown in Figure 4, which is effective for area analysis.
Additionally, while not covered here, area embedding has potential applications in downstream machine learning tasks like predicting human movement destinations and recommendations[3, 4, 31, 35, 33].
@@
The code for area analysis is available on the GitHub repository(https://github.com/UCLabNU/OpenUAS/). By downloading our data and placing it in the directory, it is possible to visualize and analyze the OpenUAS embeddings. Since the raw location data dataset for embedding cannot be shared due to privacy and confidentiality reasons, synthetic sample data is provided within the repository. By replacing these with your own dataset, it is possible to embed and compare within the same space as the OpenUAS embeddings.

### v3

@@ Abstract
We publicly release OpenUAS, a dataset of area embeddings based on urban usage patterns, including embeddings for over 1.3 million 50-meter square meshes covering a total area of 3,300 square kilometers. This dataset is valuable for analyzing area functions in fields such as market analysis, urban planning, transportation infrastructure, and infection prediction. It captures the characteristics of each area in the city, such as office districts and residential areas, by employing an area embedding technique that utilizes location information typically obtained by GPS. Numerous area embedding techniques have been proposed, and while the public release of such embedding datasets is technically feasible, it has not been realized. One reason for this is that previous methods could not embed areas from different cities and periods into the same embedding space without sharing raw location data. We address this issue by developing an anchoring method that establishes anchors within a shared embedding space. We publicly release this anchor dataset along with area embedding datasets from several periods in eight major Japanese cities.
@@
On the other hand, area embedding techniques have been proposed, aggregating features derived from location information by area and mapping them onto a vector space[29, 30, 31, 32, 33]. Such embeddings allow for advanced area analysis by aggregating information for each area, which eliminates personal data and enables searching and clustering of areas with similar functions and usage patterns. In this paper, we apply the embedding method called Area2Vec proposed by Shoji et al[34]. This method embeds areas based on stay information extracted from GPS data, resulting in embeddings called UAS (Usage of Area with Stay information) that reflect how people use the areas. The UAS can capture not only the static functions of urban areas but also the dynamic changes in area functions by tracking shifts in area usage patterns due to environmental changes(urban development, seasonal changes, pandemics). By adopting UAS as the public area embedding dataset, it becomes possible to analyze changes in area functions over multiple periods, in addition to comparing area functions between cities.
Releasing area embedding datasets provides valuable information for analyzing area functions to users who cannot directly access raw GPS datasets. Despite the demand for the public release of area embedding datasets, existing methods have not released area embedding datasets. One reason is that area embeddings, learned independently in each dataset, are unique representations. They embed the relative similarities between areas within a dataset into an embedding space. This uniqueness poses a challenge: areas from different datasets, such as cities or periods, cannot be directly compared. For example, embeddings learned independently for areas in New York and Tokyo cannot be compared due to the lack of a common embedding space. Similarly, embeddings of the same city from different periods, such as pre-COVID and post-COVID periods, are incomparable. This issue can be addressed by combining datasets of New York and Tokyo or by training new translation models[35, 36]. However, if the location data for each city is owned by different organizations, sharing the actual location data becomes difficult, making these methods impractical. Therefore, we cannot compare them with existing publicly available embeddings, even if new embeddings are generated for a different city (or a different period) using their datasets.
@@
  • The public release of area embedding data for eight major cities in Japan with anchor data based on extensive location information. (https://zenodo.org/records/13141800)
  • Development and release of tools and codes for effective use of area embedding data. (https://github.com/UCLabNU/OpenUAS/)
The next section will detail the area embedding and anchoring methods. In Data Records, we will explain the specifics of the area embedding data, and in Validation, we will assess the effectiveness and anonymity of the embedding representations and the anchoring method. Finally, in Usage Notes, we will discuss effective ways to utilize area embedding data, including the anchoring method, and provide relevant code.
@@ Methods
First, we will explain the GPS dataset we use for area embeddings and the dataset released in this paper. We release area embeddings using Area2Vec method and the anchoring method described later. Area2Vec embeds areas based on stay information in each area. We first obtained GPS data, which includes anonymized user IDs, latitude, longitude, and timestamps, from Blogwatcher Inc. This data is obtained with the use of multiple applications on mobile devices with user consent. To extract stay information from the GPS data, we used the stay estimation method by Iwata et al[37]. The resulting stay dataset includes columns for latitude, longitude, day of the week, time, and stay duration for each stay, where the day of the week and time indicate the start of the stay.
@@
The architecture of Area2Vec is shown in Figure 1[34]. The input is a one-hot vector corresponding to each area ID in the dataset, and the output is a one-hot vector corresponding to each discretized stay feature. As shown in Table 3, the discretization of stay features involves assigning the day of the week, arrival time(start time of each stay), and stay duration to discrete classes. This discretization results in the stay characteristics being classified into $2*12*7=168$ categories, and therefore, in this paper, $N^{\prime}=168$. The weight of the input layer W corresponds to the Area Embedding, and in this paper, by setting $H=8$, we obtain an 8-dimensional embedding. By learning the relationships between areas and discretized stay features, areas with similar usage patterns are embedded close to each other in the vector space, while dissimilar areas are embedded farther apart. In more detail, when a one-hot vector of each area ID is inputted, the model is trained to approximate a 168-dimensional statistical vector, which represents the aggregated discretized stay characteristics of that area. Consequently, areas with similar statistical vectors will have similar vectors, resulting in the 168 dimensions being compressed into an 8-dimensional embedding. Therefore, clustering of area embeddings yields results as shown in Figure 4. This figure will be explained in detail in the Technical Validation section.
Figure 1: Area2Vec architecture (Adapted from Shoji et al., 2021[34]).
Day of the week Weekday, Weekend or Holiday
@@
We develop an anchoring method to fix the embedding space across datasets, allowing the embedding of areas within each dataset into a common space. This enables the comparison of area embeddings across different datasets. First, we sample area data (anchor data) that includes typical areas such as residential and office districts. Then, during the embedding process, this anchor data is mixed with the dataset for training, fixing the embedding values corresponding to the Anchor. This approach fixes the latent meanings of each dimension in the embedding space across datasets. The following sections will sequentially explain the details of the anchoring method.
Figure 2: Anchor data is sampled from location dataset. Data from multiple cities and periods are embedded and clustered. Samples from each cluster are used as anchor data, and reference values are obtained as embeddings for each anchor.
First, we sample a group of anchor data $D_{a}=\{d_{a_{n}}\}$, including typical area data such as residential and office districts. The overview of the sampling of anchor data is shown in Figure 2. Each anchor data $d_{a_{n}}$ corresponds to the $n$th anchor $a_{n}$, which needs to be evenly distributed in the latent space. In other words, each anchor data $d_{a_{n}}$ should ideally represent a different area usage pattern, and $D_{a}$ as a whole should encompass a diverse range of area tendencies. To sample such $D_{a}$, we embedded a large location dataset $D^{*}$ into a single latent space, followed by clustering, and used data sampled from each cluster as the respective anchor data $d_{a_{n}}$. $D^{*}$ includes areas from eight cities in Japan, with data from Nagoya covering the years 2019-2023 (treating the same area as different for each year). The resulting $D_{a}$ includes unique area tendencies for each of the eight cities, as well as a variety of area tendencies such as pre-COVID, during the pandemic, and post-pandemic. For the public release data, we set the number of anchors $n=512$ and the amount of data for each anchor data at $20,000$. Optimization of the number of anchors and the amount of data is described in the appendix. In the $D_{a}$, timestamp and stay duration are recorded in 15-minute intervals to ensure that the stays of specific individuals cannot be extracted from $D_{a}$.
@@
The OpenUAS can be accessed at zenodo(https://zenodo.org/records/13141800). Embedding data for each dataset is aggregated every 50m or 250m mesh and stored as vector information. The total number of meshes is 1,418,827. The data columns include geocode, Latitude, Longitude, Geometry, vector, cluster5, cluster10, and cluster20. The geocode is an identifier assigned to each standard mesh, with 250m meshes having 10-digit and 50m meshes having 12-digit integer values. Latitude and Longitude indicate the center coordinates of each mesh. Geometry and Vector are lists of floating-point numbers, with the former representing the mesh polygon and the latter the vector representation of the area embedding. cluster5, cluster10, and cluster20 are the numbers assigned when the area embeddings are clustered into those respective numbers. Graphs showing the area usage patterns of each cluster in these clusterings are also included.
Additionally, we released the anchor embeddings and anchor dataset to enable users to embed their own data into the same latent space using the anchoring method. The Anchor Embeddings are defined as $512\times 8$ dimensional tensors in PyTorch and are stored as .pt files. The anchor data is a $512\times 20,000$ record .csv file containing anchor_id, arrival_time, and stay_time. All codes, including those for embedding and area analysis, are released at the GitHub repository(https://github.com/UCLabNU/OpenUAS/). Detailed explanations on how to use them are provided in the Usage Note and README.
Figure 4: Area usage patterns of each 5 cluster, including all cities and periods($D^{*}$).
@@
In this section, we demonstrate the utility of our data by practicing a part of cross-city and cross-period area analysis using clustering of area embeddings, thereby showing its effectiveness. Here, we embedded all areas from all cities and periods into the same latent space and used k-means++ for clustering. The number of clusters was set to five, but this number is arbitrary and can be increased for more detailed analysis. Graphs for when the number of clusters was set to 10 and 20 are included in the appendix section. For different periods of data in Nagoya, we treated each period as a distinct area. For example, the areas of Nagoya in January 2021 and February 2021 are considered as different areas. This allows us to analyze how area clusters change over time. The resulting stacked graph is shown in Figure 4. The horizontal axis represents time, and the vertical axis represents the number of visits per area every 30 minutes, with the color indicating the length of stay. On the right side of the graph, we display labeled interpretations(manually predicted) based on the tendencies of each cluster, along with the count of areas belonging to each cluster. For instance, the first and second clusters, with many short stays on weekends compared to weekdays, can be inferred as areas used like entertainment districts. The third cluster, with many stays over 360 minutes during weekdays, appears to be an office district cluster, while the fourth cluster, with frequent stays under 30 minutes during weekday mornings and evenings, seems to be areas with stations or streets. The fifth cluster, with many stays over 720 minutes, mainly during weekend nights, is a residential area cluster. Thus, areas can be clustered based on their functional differences, and each interpretation of the cluster can be labeled based on the inferred stay tendencies from the graphs.
Next, we visualize the geographical map of these area clusters. We begin by showing the map of area clusters for each city in April 2021 in Figure 5. From the figure, we can observe the distribution of areas with distinct usage patterns, such as central train stations and street areas, industrial and office regions in downtown and bay areas, and suburban residential districts. An important point to note here is that despite each area being embedded without sharing location datasets, the anchoring method ensures that each color representing an area function has the same meaning across different cities. Simultaneously, other researchers can embed and analyze areas in other cities using their own data. Thus, our data and method facilitate cross-city analysis of area functions.
@@
Next, we evaluate cross-period area analysis. We present the distribution of area clusters in the central part of Nagoya for April from 2019 to 2023 in Figure 6. Area (a) around Nagoya Station is a region where commercial facilities, restaurants, and offices are located. In 2019, many of these areas belonged to the entertainment and restaurant cluster, but in 2020, many shifted to the workplace cluster. This change reflects the impact of the COVID-19 pandemic in 2020, indicating a decrease in visitors to entertainment and restaurant areas due to self-quarantines. Area (b) is Sakae, a major entertainment area in Nagoya. Here, too, the function of entertainment areas decreased in 2020, but there is a visible recovery of this function in 2021. Area (c) is around Nagoya Castle, a tourist spot, with government buildings such as the city hall to its south. In 2020, this area transitioned to a station and street area, likely due to decreased visitors for sightseeing purposes, while visits for commuting to the workplace area south of Nagoya Castle remained. Annual data, as illustrated here, enable analysis of area changes before and after the COVID-19 pandemic.
@@
Finally, we visualize the monthly changes of areas within the same region. We have chosen and visualized the three months from September to November 2021 from the entire year, as shown in Figure7. Area (d) includes a park, where a festival event was held from September 26 to October 31, 2021. The impact of this event, coupled with the lifting of the state of emergency from August 27 to September 30, 2021, can be seen in the changing functions of the area. In area (e), a large shopping mall opened on October 27, 2021. This led to the rise of an entertainment area in a region that previously only had workplaces and residential areas. Thus, using annual and monthly data, it is possible to analyze changes in the functions of the same area over time.
@@
Validation of anchoring method
Next, we validate the effectiveness of the anchoring method. The key aspect of the anchoring method is the ability to embed areas in different datasets into the common embedding space without sharing datasets. In traditional methods without anchoring technique, even if the exact same dataset is embedded, the first and second embeddings are positioned far apart, failing to share the embedding space. In this method, areas in each dataset are embedded into a common embedding space $E^{+}$ through the anchor data $D_{a}$. To evaluate whether they are appropriately embedded in the common space, we embed $D^{+}$ twice, comparing the results with and without the anchoring method. If the anchoring is not functioning correctly, the first and second embeddings will be significantly distant from each other in the latent space. We then define Misalignment, as shown in the following equation, as the discrepancy between the obtained embedding $E$ and $E^{+}$
$M\!isalignment(E,E^{+})=\frac{1}{|M|}\sum\limits_{m\in M}D\!istance(e_{m},e^{+}_{m})$ (3)
$M$ represents the set of all areas in the dataset $D^{*}$, with each area denoted as $m$ and the total number of areas in $M$ as $|M|$. $e_{m}$ represents the retrained embedding corresponding to $m$, whereas $e_{m}^{+}$ represents the embedding for $m$ in $E^{+}$. A smaller $Misalignment(E,E^{+})$ indicates a greater success in embedding each area into the same latent space as $E^{+}$ in the retraining. $Distance$ is a distance function for which both Euclidean distance and Cosine distance are used. Cosine distance is calculated as 1 minus the cosine similarity. Here, we compared the original Area2Vec without anchoring and Area2Vec with anchoring. Results are shown in Figure 8. From these results, it is evident that the original method yields different embeddings for independent learning, indicating that it does not embed in a common latent space. In contrast, our proposed anchoring method resulted in a smaller variance in embeddings for the same data, demonstrating the ability to embed areas into the same representation space.
@@
Simultaneously, it is necessary to evaluate whether the ability in capturing area usage patterns has been compromised by anchoring. This concern arises because the anchoring method may distort the update of weights during learning via anchors, potentially preventing the acquisition of optimal area embeddings for the training data. In particular, area data with usage patterns not included in anchor data may not be representable in the embedding space of OpenUAS. To validate this, we conducted the following experiments for each dataset comprising $D^{*}$.
@@
If the performance using anchoring significantly declines compared to when learning individually, it may indicate that the anchor data does not fully capture the latent patterns of areas within the target data, failing to anchor them appropriately in the ideal space.
Although area embedding is an unsupervised learning process, poses challenges for direct performance validation. To address this, we introduce the metric $ApproximationLoss$. Area embedding serves as a method of dimension reduction. During the training process, the product of the embedding $e_{m}$ for area m and the output matrix $W^{\prime}$ is optimized to better approximate the frequency $f_{m}$ of the discretized stay feature for each area. Here, $f_{m}$ is a 168-dimensional vector representing the frequency of the one-hot for discretized stay features. Essentially, the performance of area embedding improves as the product of $e_{m}$ and $W^{\prime}$ more closely approximates $f_{m}$. This discrepancy is quantified by $ApproximationLoss$. The corresponding formula is presented below.
@@
Here, $f_{m}$ represents the true frequency of discretized stay features in area $m$, and $\hat{f_{m}}$ represents the approximate frequency calculated as the product of $e_{m}$ and $W^{\prime}$. We conducted the previously described experiments on $D_{tokyo}$ to $D_{2023}$, which make up $D^{*}$, calculating and comparing the $ApproximationLoss$ learned using the original Area2Vec(without anchoring) and the anchoring method. The results are shown in Figure 9. The results reveal that there is no significant difference in the $ApproximationLoss$ values before and after the application of the anchoring method. However, it is also noted that the $ApproximationLoss$ values differ by a few points regardless of the application of the anchoring method, with the differences being especially prominent across different years. This variation is believed to be caused by the differences in latent trends of areas across the datasets. For instance, the year 2019 shows worse metrics compared to other datasets, which could be attributed to the greater diversity in area characteristics, making it more challenging to represent them with fewer embedding dimensions. The variations in embedding accuracy due to differences in cities and years suggest that further evaluation is needed using data from cities outside Japan. Nonetheless, the minor deterioration in embedding accuracy before and after applying anchoring indicates that the method is effective.
@@
In the second step, these statistical vectors are embedded into an 8-dimensional latent space, further abstracting the information. As shown in the Technical Validation section, the degree of match between the approximated and actual statistical vectors for each mesh is calculated by $ApproximationLoss$. Therefore, data users can only know the approximate values of the statistics for each mesh. Moreover, $ApproximationLoss$ tends to be higher for meshes with more user counts and lower for those with fewer. Figure 10 illustrates the relationship between user count per mesh and $ApproximationLoss$. From Figure 10, it is apparent that in meshes with 10-20 user counts, the values are higher and decrease as the user count increases. This result suggests it is particularly difficult to infer related statistical information in meshes with fewer users. Therefore, our data does not contain personal data and only provides approximate values for any statistical information that could indirectly estimate personal data. From these aspects, it can be concluded that our data is sufficiently anonymized.
@@
This involves clustering a group of target area embeddings into any number of clusters, forming clusters of areas with similar functions. This allows for the classification and analysis of urban sections based on their functions and usage. Our code includes the tool for clustering area embeddings into any number of clusters and the tool for geographical visualization.
@@
As explained in the Technical Validation, area trends can be predicted through the approximation of discretized stay statistics. While not perfectly matching actual statistics, this method can reproduce stay trend graphs as shown in Figure 4, which is effective for area analysis.
Additionally, while not covered here, area embedding has potential applications in downstream machine learning tasks like predicting human movement destinations and recommendations[3, 4, 31, 35, 33, 40].
@@
The code for area analysis is available on the GitHub repository(https://github.com/UCLabNU/OpenUAS/). By downloading our data and placing it in the directory, it is possible to visualize and analyze the OpenUAS embeddings. Since the raw location dataset for embedding cannot be shared due to privacy and confidentiality reasons, synthetic sample data is provided within the repository. By replacing these with your own dataset, it is possible to embed and compare within the same space as the OpenUAS embeddings.